Circular No. 6731

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

USNO 1425.09823278 J.-y. Hu, Y.-l. Qiu, W.-d. Li, and J.-y. Wei, Beijing Astronomical Ob-

servatory (BAO); and A. Esamdin, Urmqi Astronomical Station, report their discovery of a new apparent cataclysmic variable star: “The discovery was made during the BAO supernova survey on Aug. 28, using the 0.60-m telescope at Xinglong Station. The R magnitudes of this star were 13.3 and 13.5 on Aug. 28 and 31, respectively. The USNO A1.0 catalogue lists the R and B magnitudes of a star in this location as 19.9 and 20.3, respectively, with the position given as α = 19^{h}27^{m}11^{s}.63, δ = +54^{o}17′51′′.5 (equinox 2000.0). A spectrogram obtained with the 2.16-m telescope at Xinglong Station on Aug. 31 shows typical features of cataclysmic variables during outburst, with absorption Balmer lines and weak He I lines (including a feature at 447.1 nm).”

GRB 970828 K. Yanagisawa and Y. Nakada, Kiso Observatory, University of Tokyo;

and M. Abe and Y. Ishibashi, Institute of Space and Astronautical Science, report that a search was made with the Kiso Schmidt telescope on Aug. 29.46–29.52 and 30.45–30.63 UT for an optical counterpart to GRB 970828, around the position given by the ASCA team on IAUC 6729. Three 15min I-band CCD images (size 12′.5 × 12′.5) revealed no new objects brighter than the limiting magnitude of the second Palomar Observatory Sky Survey. However, two faint objects were seen very close to the x-ray position; the closer object (I ∼ 20.7) is at α = 18^{h}08^{m}29^{s}.48, δ = +59^{o}19′24′′.3 (equinox 2000.0), while a second object (I = 20.5) has position end figures 29^{s}.48, 35′′.1. The positional uncertainties are 0^{s}.03 in α and 0′′.2 in δ. No light variation > 0.5 mag was detected during the two-day observations.

SUPERNOVA 1997dd IN NGC 6060 Reported unfiltered CCD magnitudes (GSC references stars): Aug.

29.89 UT, 16.7 (S. Moretti and S. Tomaselli, Forlí, Italy); 29.905, 17.1 (L. Boschini, Bergamo, Italy).

SUPERNOVA 1997cx IN NGC 3057 Unfiltered CCD magnitude estimate by Moretti and Tomaselli: Aug.

29.9 UT, 16.7.

1997 September 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green